Buffer solutions:


1. Calculate the pH of a carbonate buffer system that contains 0.0065 M HCO3 and 0.0084 M CO32-. The Ka of HCO3- is 4.7 X 10-11.


1
Expert's answer
2021-04-19T05:08:42-0400

HCO3- = H+ + CO32-


Ka=[H+]*[CO32-]/[HCO3-] = [H+]*c(CO32-)/c(HCO3-) = 4.7*10-11


[H+] = Ka*c(HCO3-)/c(CO32-) = 4.7*10-11*0.0065/0.0084 = 3.6369*10-11


pH = -log10([H+]) = -log10(3.6369*10-11) = 10.44